OpenStudy (anonymous):

Gloria is 20 years older than Reggie. Five yeas ago she was five times as old as he was. How old is each now?

OpenStudy (diyadiya):

Let Reggie be x years old Gloria's age = x+20 5 years ago Gloria was five times as old as Reggie (x+20)-5=5(x-5) x+15=5x-25 25+15=5x-x 40=4x x=10 Reggie is 10yrs old Gloria is 10+20=30 yrs old

OpenStudy (diyadiya):

Did you understand?

OpenStudy (anonymous):

Yes, wow. Thank you! You're amazing!

OpenStudy (diyadiya):

Anytime =)